OpenStudy (anonymous):

Evaluate the following limit or explain why they do not exist. lim as theta approaces pi/2^- tantheta^cos(thetha)

OpenStudy (anonymous):

I'm not sure if i'm doing this correctly, the problem is: \[\lim_{\theta \rightarrow \pi/2^-} (\tan \Theta)^{\cos \Theta}\] USing this formula \[b^{x}= e^{xlnb} \] I get \[\cos \Theta \ln(\tan \Theta)\] According to my steps for L Hopital: 1. Find indeterminate form 2. Get ratio 3. get derivative, no quotient rule used for the ratio 4. Find the limit 5. repeat if in indeterminate form Ratio: \[\ln(\tan \Theta)/\sec \Theta = f \prime/ g \prime = \frac {\sec^2 \Theta/\tan \Theta}{\sec \Theta \tan \Theta}\] Is this right so far?

OpenStudy (anonymous):

*that should be ratio and the derivative

OpenStudy (anonymous):

I also had a question about why is it that when I evaluate: \[\sec^2 \Theta \] as Theta approaches pi/2^- the answer is infinity, I thought it would be 0 since costheta at pi/2^- is zero.

myininaya (myininaya):

sec and cos are reciprocals of one another

myininaya (myininaya):

cos =0 then sec=1/0 (or does not exist)

myininaya (myininaya):

and yes you are right so far do another thing of l'hospital

myininaya (myininaya):

before you do it make sure you clean up that nasty fraction.